OpenStudy (thefluffmuncher):

Write the equation in standard form: 6 – 2(x – y) = 3x + 5

OpenStudy (thefluffmuncher):

5x – 2y = 1 why is this the answer i'm confused

OpenStudy (thefluffmuncher):

@dan815

OpenStudy (thefluffmuncher):

@Destinymasha

OpenStudy (godlovesme):

can u isolate the variable y ? @thefluffmuncher

OpenStudy (snowsurf):

You basically simplify the expression into standard form.

OpenStudy (thefluffmuncher):

idk i just distributed to get rid of the parenthesis

OpenStudy (thefluffmuncher):

i know the formula Ax+By=c

OpenStudy (snowsurf):

6 – 2(x – y) = 3x + 5 distribute first... 6 -2x + 2y = 3x + 5 Combine like terms -5x + 2y = -1 And they multiply by -1 to get 5x -2y = 1

OpenStudy (thefluffmuncher):

Why did you multiply the whole equation by -1?

OpenStudy (snowsurf):

They have it in standard form because if you are doing data analysis its easy to see the slop and the x and y intercepts. As for multiplying by -1 do you have to do it? No. I could had left it as -5x + 2y = -1. I did that to match your answer you posted.

OpenStudy (thefluffmuncher):

oh ok

OpenStudy (thefluffmuncher):

thanks